OpenStudy (anonymous):

Solve for x 10^(log12x+6)=5

OpenStudy (asnaseer):

do you know what this equals:\[b^{\log_bx}=?\]

OpenStudy (anonymous):

no i dont

OpenStudy (asnaseer):

ok, look at it this way, you this rule I presume - if:\[\log_bx=y\implies x=b^y\]

OpenStudy (asnaseer):

*you know this...

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

good, so now substitute into the last expression the value for y and we get:\[x=b^y=b^{\log_bx}\]

OpenStudy (asnaseer):

since \(y=\log_bx\)

OpenStudy (asnaseer):

make use of this rule to simplify your equation

OpenStudy (asnaseer):

I assume your equation is:\[10^{\log_{10}(12x+6)}=5\]

OpenStudy (anonymous):

yea so is it 10^5=12x+6?

OpenStudy (asnaseer):

no - why did you put \(10^5\)?

OpenStudy (asnaseer):

the 5 should just remain as it was

OpenStudy (asnaseer):

\[10^{\log_{10}(y)}=y\]

OpenStudy (asnaseer):

that is the identity you need to use here.

OpenStudy (anonymous):

ahh ok i get it. thank you so much

OpenStudy (asnaseer):

yw :)